The mwan numbers of cars passing at a junction is 50. The passage of cars follow a Poisson distribution. Find the probability of less than 40 cars passing.


1
Expert's answer
2022-03-15T19:42:00-0400

"P(X=x)=\frac{e^{-\lambda}\lambda^x}{x!},x=0,1,2,3..."

In this case "P(X<40)=P(X=0)+P(X=1)+...+P(X=38)+\nP(X=39)"

"=0.06457"